Please help me solve this problem, I do not understand this. A rectangle has a length of x + 2, a width of 7, and a perimeter of 32 what is the value for x?
Answer by edjones(7569)

2L+2w=P
2(x+2)+2*7=32
2x+4+14=32
2x+18=32
2x=14
x=7
.
Ed